Hey,

i need help with problem: how many pounds of candy worth $3.60 per pound must be added to 10 pounds of chocolate worth $2.40 per pound to form a mixture worth $3.30 per pound?

thanks

Answer by stanbon(48510)

i need help with problem: how many pounds of candy worth $3.60 per pound must be added to 10 pounds of chocolate worth $2.40 per pound to form a mixture worth $3.30 per pound?
------------------
$3.60 Candy DATA:
Amount = x lbs ; Value = 3.60x dollars
--------------------
$2.40 Candy DATA:
Amount = 10 lbs; Value = 24 dollars
--------------------
Mixture DATA:
Amount = x+10 lbs ; Value = 3.30(x+10) = 3.3x + 33
--------------------
EQUATION:
value + value = value
3.6x + 24 = 3.3x + 33
0.3x = 9
x = 30 lbs (amount of $3.60 Candy tht must be added)
=============================
Cheers,
Stan H.